September 10

 

Machine preparation for Wednesday

 

Final tests finished around midnight Tuesday

 

Check SPS OK to TEDs overnight.

Pull it all together from 06h.

       Injectors.

       LHC interlocks.

       LHC machine.

       LHC settings.

       Experiments and associated interlocks.

 

Status 06.30

Lost the cryo in 78 due to turbine stop at 04.00. Recovering.

Also search down in point 2. Patrol needed.

 

 

Patrol  done by 08h. Getting SPS ready.

08.30 Beam to TEDS both lines. Interleaved injection.

09.00 Switch to beam 1 only. Kickers executing soft start.

09.30 Beam to TDI. Kickers on.

09.40 TDI out. Beam to collimators point 3.

09.45 Collimators IP3 out. Beam to left of point 5.

09.55 Collimators point 5 out, beam ~to point 6 but not to the extraction line.

10.00 Beam to dump block (steered with DC bump). Half way round ring1!

10.08 Beam to collimators in point 7.

10.12 Beam to point 8 collimators.

 

10.30 Beam round > 1 turn  image on screen showing beam on first and second turns  trajectory

With a bit of correction - beam makes 3 turns (screen image)

 

High spot of the day is to have made Google

 

lhc google

 

Images from Atlas  from beam on the collimator in front of them Atlas 1  Atlas 2  (& thanks for the champagne)

and Alice's pixel dectector

 

Non-serious cryogenics problem delays injection of beam 2 until 13:30

 

Switch to beam 2 only.

13.30 onto the TDI

13.40 TDI out, beam to point 3.

Cryogenic instability needs a little more time

13.55 Beam to 6, steered into start of dump line

14.05 Beam to left of 6. Lost. Investigating.

14.25 Problem understood. Beam to right of 5. 20 shots for CMS.

14.40 Beam round to 3 (after one correction in 4). Few corrections needed.

14.45 Beam to ALICE.

14.50 Beam to ATLAS. Few shots for them.

Few minutes no beam. Reboot for screen server.

15.00 Beam > 1 turn

 

We take a breather - back soon for some beam based measurements.

 

16.00 Beam 2 to collimators ATLAS

Dispersion data

Kick response data

In parallel BI looking at BPM acquisition for >1 turn

 

18.00 Inject and dump commissioning (all collimators out)

Inject and dump up to 9 turns OK.

20.00 BPM acquisition for multiple turns sorted out

Trimming Qh, Qv, Q'

21.30 Beam 2 makes at least 300 turns!

Fast BCT data and tune measurements - Qh .31 Qv .23

23.00 Start systematic polarity checks of orbit correctors

03.00 RB.34 trip, followed by problems with a number of circuits in 56, D1.R8 and QT11.R2

08.00 Access in 4 points for above problems